A random sample of enrollments from medical schools that specialize in research and from those that are noted for primary care is listed. Find the 90% confidence interval for the difference in the means.
Research

474 571 596 658 782 477 667 414 807 437
571 688 689 692 280 414 874

Primary care

783 609 420 736 556 469 371 115 442 597
291 278 670 556 522 310

  

18.5799 < μ1 - μ2 < 202.9715

  

-9.7075 < μ1 - μ2 < 231.2589

  

24.8659 < μ1 - μ2 < 196.6855

  

6.0077 < μ1 - μ2 < 215.5437
Answer #1

Ans:

n x-bar s
Research 17 593.5882 160.7008
Primary care 16 482.8125 181.2292

Assumption:Population variances are not equal.

Point estimate=593.5882-482.8125=110.7757

df=16-1=15

t*=tinv(0.1,15)=1.753

Margin of error=1.753*SQRT((160.7008^2/17)+(181.2292^2/16))=104.7680

90% confidence interval for difference in means

=110.7757+/-104.768

=(6.0077, 215.5437)

Last option is correct.